### v1

# Log-concave functions and transformations thereof

Abstract
I summarize BagBerg05Logconcave’s review on log-concave functions, make several corrections, and augment the discussion with further results that can be useful in obtaining monotone hazard rate. I also provide an application of monopoly pricing.
1 Property of Log-Concave Functions
Throughout, we are interested in a twice continuously differentiable function $f$ mapping from a connected set $D\subseteq\mathbb{R}$ to $\mathbb{R}_{++}$.
Definition 1.
$f(x)$ is log-concave on an interval $(a,b)\subset D$ if
$(\log(f(x)))^{\prime\prime}\leq 0,\forall x\in(a,b).$
It is strictly log-concave on $(a,b)$ if the above inequality is strict.
Remark. $f$ is said to be a log-concave function if $f$ is log-concave on $D$.
The following propositions describe some properties of log-concave functions. The first is immediate if one write the derivative condition in Definition 1 explicitly.
Proposition 1.
The followings are equivalent:
  • 1.
    $f(x)$ is log-concave on $(a,b)$;
  • 2.
    $\frac{f^{\prime}(x)}{f(x)}$ is decreasing on $(a,b)$;
  • 3.
    $f^{\prime\prime}\cdot f-(f^{\prime})^{2}\leq 0$ on $(a,b)$.
Proposition 2.
The followings are true:
  • 1.
    A nonnegative concave function is log-concave.
  • 2.
    A log-concave function is quasi-concave.
See Appendix A.1. ∎
Proposition 3.
If $f$ is a log-concave function, then $f$ is unimodal.
We are to show that if $f$ is log-concave, there is at most a single peak or a plateau. In other words, the solution to problem $\max_{x\in D}f(x)$ constitutes at most one connected interval. The first-order necessary condition is $f^{\prime}(x)=0$ for interior local maximum. Since $f^{\prime}(x)/f(x)$ is decreasing, the equation $f^{\prime}(x)/f(x)=0$ has solutions in at most one connected set. Since $f(x)>0$, this is equivalent to $f^{\prime}(x)=0$ having solutions in at most one connected set. ∎
Definition 2 (Monotone likelihood ratio property).
A family of probability density functions $\{f(x|\theta)\}$ has the monotone likelihood ratio property (MLRP) if, for any $\theta_{1}<\theta_{2}$, the likelihood ratio $\frac{f(x|\theta_{2})}{f(x|\theta_{1})}$ is non-decreasing in $x$.
Proposition 4.
A family of probability density function $\{f(x-\theta),\theta\in\mathbb{R}\}$ satisfies MLRP if and only if $f$ is log-concave.
See Appendix A.2. ∎
Remark. For example, the family of normal distributions $\{N(\mu,\sigma^{2})\}$ is commonly used by varying $\mu$. It satisfies the MLRP, and each normal density function is log-concave. In fact, one can use any log-concave density function to construct a family that satisfies MLRP, according to Proposition 4.
Theorem 1 (BagBerg05Logconcave).
If $f(x)$ is log-concave on $(a,b)$, then
  • 1.
    $F(x):=\int_{a}^{x}f(t)\,dt$ is strictly log-concave on $(a,b)$;
  • 2.
    $\overline{F}(x):=\int_{x}^{b}f(t)\,dt$ is strictly log-concave on $(a,b)$.
We begin by proving $F(x)$ is log-concave. By Proposition 1, it suffices to show that $F^{\prime\prime}\cdot F-(F^{\prime})^{2}\leq 0$ on $(a,b)$, or
$f^{\prime}\cdot F-f^{2}\leq 0$ (1)
on $(a,b)$. Since $f$ is log-concave on $(a,b)$, $f^{\prime}(x)/f(x)$ is decreasing on $(a,b)$. Then,
$\frac{f^{\prime}(x)\cdot F(x)}{f(x)}$ $=\frac{f^{\prime}(x)}{f(x)}\int_{a}^{x}f(t)\,dt$
$\leq\int_{a}^{x}\frac{f^{\prime}(t)}{f(t)}f(t)\,dt$
$=\int_{a}^{x}f^{\prime}(t)\,dt=f(x)-f(a),$
i.e., $\frac{f^{\prime}(x)\cdot F(x)}{f(x)}\leq f(x)-f(a)$. Rearranging yields
$f^{\prime}\cdot F-f^{2}\leq-f(a)\cdot f<0,$
implying inequality (1) holds. Hence, $F(x)$ is strictly log-concave on $(a,b)$.
On the other hand, $\overline{F}(x)$ being log-concave requires
$f^{\prime}\cdot\overline{F}-(f)^{2}\leq 0.$
Using the same logic,
$\frac{f^{\prime}(x)\cdot\overline{F}(x)}{f(x)}$ $=\frac{f^{\prime}(x)}{f(x)}\int_{x}^{b}f(t)\,dt$
$\geq\int_{x}^{b}\frac{f^{\prime}(t)}{f(t)}f(t)\,dt$
$=\int_{x}^{b}f^{\prime}(t)\,dt=f(b)-f(x),$
which implies
$f^{\prime}\cdot\overline{F}-(f)^{2}\leq-f(b)\cdot f(x)<0.$
Therefore, $\overline{F}(x)$ is strictly log-concave on $(a,b)$. ∎
Corollary 1.
The reliability function $H(x)=\int_{x}^{b}\overline{F}(t)\,dt$ is log-concave on $(a,b)$ if $\overline{F}(x)$ is log-concave on $(a,b)$.
Corollary 2.
The mean residual lifetime function $\frac{H(x)}{H^{\prime}(x)}=\frac{\int_{x}^{b}\overline{F}(t)\,dt}{\overline{F}(x)}$ is decreasing on $(a,b)$ if $\overline{F}(x)$ is log-concave on $(a,b)$.
The corollary immediately follows from Theorem 1. Note that a stronger sufficient condition for both corollaries to hold is $f(x)$, as opposed to $\overline{F}(x)$, being log-concave on $(a,b)$.
2 Other Transformations
In the previous section, I have shown that log-concavity is preserved under integration. I show other transformations that preserves log-concavity in this section.
2.1 Multiplication
Theorem 2.
If $f$ and $g$ are log-concave, then $h(x)=f(x)\cdot g(x)$ is log-concave.
Since concavity is preserved under sums, $\log h(x)=\log f(x)+\log g(x)$ is concave in $x$, i.e., $h$ is log-concave. ∎
Remark. Summation does not preserve log-concavity in general. To gain intuition, consider two unimodal density functions with different modes. Their sum may not be unimodal anymore.
2.2 Composition
Theorem 3.
Suppose $f$ is log-concave on $(a,b)$. Let $t$ be a monotonic, twice-differentiable function mapping from $(a^{\prime},b^{\prime})$ to $(a,b)=(t(a^{\prime}),t(b^{\prime}))$. Define the function $f_{t}$ with support $(a^{\prime},b^{\prime})$ so that for all $x\in(a^{\prime},b^{\prime})$, $f_{t}(x)=f(t(x))$.
  • (i)
    If $f$ is increasing and $t$ is concave, then $f_{t}$ is log-concave. [^1] BagBerg05Logconcave implicitly assume $f$ is increasing in their related results (Theorem 7).
  • (ii)
    If $f$ is decreasing and $t$ is convex, then $f_{t}$ is log-concave.
Examining the second-order derivative on $f_{t}$ with respect to $x$, we have
$\frac{(t^{\prime})^{2}[f^{\prime\prime}\cdot f-(f^{\prime})^{2}]+f\cdot f^{\prime}\cdot t^{\prime\prime}}{f^{2}}.$
A sufficient condition for the above expression to be nonpositive (so that $f_{t}$ is log-concave) is that $f\cdot f^{\prime}\cdot t^{\prime\prime}\leq 0$, because the first term in the numerator is nonpositive by log-concavity of $f$ (c.f. Proposition 1). Therefore, either (i) or (ii) is a sufficient condition for $f_{t}$ to be log-concave. ∎
Since linear transformation is both concave and convex, the following corollary is immediate.
Corollary 3.
Suppose $f$ is log-concave on $(a,b)$. Let $t$ be a linear transformation from $\mathbb{R}$ to $\mathbb{R}$. Define a function $f_{t}$ with support $(t(a),t(b))$, then $f_{t}$ is log-concave on $(a,b)$.
It is sometimes useful to set the transformation $t$ to be the ratio $\Gamma(x):=\frac{F(x)}{f(x)}$. It is easy to verify that for uniform distributions, $\Gamma(x)$ is linear. For exponential distributions, $\Gamma(x)$ is convex, so $f\circ\Gamma$ is log-concave if $f$ is log-concave and decreasing. The following proposition shows, which is less obvious, that $\Gamma(x)$ is also convex for normal distributions. [^2] To avoid algebraic burden, the proposition only argues for standard normal distribution.
Proposition 5.
For the standard normal distribution with density function $\varphi(x)$ and distribution function $\Phi(x)$. Then $\Gamma(x):=\frac{\Phi(x)}{\varphi(x)}$ is convex.
See Appendix A.3. ∎
2.3 Truncation
Theorem 4 (BagBerg05Logconcave, Theorem 9).
If a probability distribution has a log-concave density function (resp. cumulative distribution function), then any truncation of this probability distribution will also have a log-concave density function (resp. cumulative distribution function).
The proof is omitted for it is verbatim in BagBerg05Logconcave. Nevertheless, I discuss the example of truncated normal distributions. [^3] See Section 6 of BagBerg05Logconcave for a nice and thorough discussion on examples of log-concave distributions.
Truncated Normal Distributions.
Denote the truncated normal distribution by $\texttt{trunc}_{[a,b]}N(\mu,\sigma^{2})$ for shorthand, meaning that the distribution is $N(\mu,\sigma^{2})$ conditional on only assumes values in the interval $[a,b]$. Then, $\texttt{trunc}_{[a,b]}N(\mu,\sigma^{2})$ has a conditional cumulative distribution function defined as [^4] See, e.g., Greene2002, pp.757-759.
$F(x)=\frac{\Phi(\xi)-\Phi(\alpha)}{\Phi(\beta)-\Phi(\alpha)},\;x\in[a,b]$
with $\Phi(\cdot)$ being the cdf of the standard normal distribution; $\alpha:=\frac{a-\mu}{\sigma}$ is the standardized lower bound; $\beta:=\frac{b-\mu}{\sigma}$ is the standardized upper bound; $\xi:=\frac{x-\mu}{\sigma}$ is the standardized statistic. The probability density function is
$f(x)=\frac{\varphi(\xi)}{\sigma(\Phi(\beta)-\Phi(\alpha))},\;x\in[a,b]$
Therefore, $\texttt{trunc}_{[0,1]}N(\mu,\sigma^{2})$ has a cdf of
$F(x;\mu,\sigma)=\frac{\Phi(\frac{x-\mu}{\sigma})-(1-\Phi(\frac{\mu}{\sigma}))}{\Phi(\frac{1-\mu}{\sigma})+\Phi(\frac{\mu}{\sigma})-1},\;x\in[0,1]$
and a pdf of
$f(x;\mu,\sigma)=\frac{\varphi(\frac{x-\mu}{\sigma})}{\sigma[\Phi(\frac{1-\mu}{\sigma})+\Phi(\frac{\mu}{\sigma})-1]},\;x\in[0,1]$
The truncated normal distribution preserves the log-concavity.
Uniform Limit as $\sigma\rightarrow\infty$, $\mu=0.5$. I claim that $F(x;0.5,\sigma)$ and $f(x;0.5,\sigma)$ converges to uniform distribution function and density function, respectively, as $\sigma$ goes to infinity. By direct calculation,
[…]
and
[…]
3 Application: Log-Concave Demand Curve
In this section, I provide an application of monopoly pricing to illustrate a role of log-concave distributions play in economic analysis.
There is a single product market. Consider a continuum of consumers, each of whom has one unit of indivisible demand of the product. Each consumer’s value of the unit of demand is $v\in[0,1]$ drawn i.i.d. from a distribution $G$ with strictly log-concave density gg. [^5] A weaker condition is to assume 1G(v)1-G(v) is strictly log-concave, with which all discussion in this section goes through.
There is a monopoly firm serving the product. [^6] One could immediately tell from a mechanism design perspective, that $g$ or 1G1-G being log-concave implies the virtual value $\psi(v):=v-\frac{1-G(v)}{g(v)}$ is increasing. Then, the revenue maximizing mechanism is a posted price mechanism with $p=\psi^{-1}(c)$ (see, e.g., borgers2015). In this section, I deliberately use the more classical demand curve approach to illustrate the relationship between log-concave value distribution and the property on the demand curve (BulowRoberts89 first formally point out the link between the two). The firm posts a unit price $p\geq 0$. Then the market demand is $q(p)=\Pr(v\geq p)=1-G(p)$. The monopolist solves
$\max_{p}(p-c)q(p),$
where $c\geq 0$ is the firm’s marginal cost of production. Assume no fixed cost. The following proposition says strict log-concavity of the demand function is a stronger property than the revenue being strictly concave. [^7] More commonly seen in the IO and regulation literature. See, e.g., Assumption 1 and Proposition 2 of WeiZouRegulationNoSubsisdy. Moreover, the optimal price has a decreasing markup as cost goes up.
Proposition 6.
The revenue function of the monopoly problem is strictly concave (w.r.t. quantity $q$) . Moreover, the monopoly price is uniquely characterized as a solution to
$p=c+\frac{1-G(p)}{g(p)}$
with the markup $\frac{1-G(p)}{g(p)}$ strictly decreasing as $c$ increases.
First, we prove that the monopolist’s revenue is strictly concave in $q$. Strict log-concavity implies $G$ is strictly increasing, which means its inverse is well-defined. Rewriting the monopolist’s problem by substituting $p(q)=G^{-1}(1-q)$, we get
$\max_{q}[G^{-1}(1-q)-c]q$
Taking derivative w.r.t. $q$, we get
$\underbrace{p-\frac{1-G(p)}{g(p)}}_{\text{marginal revenue}}\quad-\quad c.$
Note that the marginal revenue is strictly increasing in $p$ because the term $\frac{1-G(p)}{g(p)}$ is decreasing by log-concavity of $g$. Meanwhile, $p$ is strictly decreasing with $q$, which implies the marginal revenue $p(q)-\frac{1-G(p(q))}{g(p(q))}$ is strictly decreasing in $q$. That is, the revenue function is strictly concave in $q$.
Since the revenue function is strictly concave, the first-order condition (marginal revenue equal to marginal cost) is sufficient and necessary for the unique interior solution, i.e.,
$p(c)=c+\frac{1-G(p(c))}{g(p(c))}.$
One can reinterpret the term $\frac{1-G(p)}{g(p)}$ as the monopoly markup. Standard implicit function theorem shows $p$ increases as $c$ increases, which implies the markup $\frac{1-G(p)}{g(p)}$ decreases in $c$. ∎
[…]
Figure 1: Illustration of decreasing markup
Remark. The fact that the price markup decreases in cost $c$ is a distinctive feature of log-concave demand. This property does not hold for general demand systems with concave revenue; for example, constant-elasticity demand exhibits increasing markup.
The intuition is most transparent from the perspective of information rents. For a log-concave distribution $G$, the information rent of type $v$, given by
$\frac{1-G(v)}{g(v)},$
is strictly decreasing in $v$. Serving any type incurs incentive externalities on all higher types, so the marginal benefit of serving consumers rises more steeply than in the first best. As $v$ approaches its upper bound, these incentive effects disappear—hence “no distortion at the top” (see Figure 1).
Under the first-best benchmark, the planner serves consumers until their value equals cost: $\hat{v}=c$. A monopolist instead sets output by equating marginal revenue and cost:
$\text{MR}(\hat{v})=\hat{v}-\frac{1-G(\hat{v})}{g(\hat{v})}=c,$
leading to under-provision. Implementing this through a unit price requires
$p-\frac{1-G(p)}{g(p)}=c,$
so the indifferent consumer has value $\hat{v}=p$. Rearranging gives
$p(c)=c+\frac{1-G(\hat{v})}{g(\hat{v})},\hat{v}=p(c),$
revealing that the price markup equals the information rent of the marginal (indifferent) consumer. Because the monopolist internalizes the incentive externality of extending trade, its pricing rule embeds exactly this rent—capturing the misalignment between monopolist and planner.
As cost $c$ increases, the indifferent consumer must lie at a higher value. Since information rents decrease in $v$, the markup necessarily falls. Intuitively, the monopolist’s objective becomes more aligned with the social planner when serving higher-valued consumers, explaining why markups decrease with cost.
Demand-side interpretation. The same phenomenon can be understood directly from the shape of the demand curve. For log-concave $G$, the demand curve q(p)=1G(p)q(p)=1-G(p) becomes more elastic as price increases:
Proposition 7.
The demand curve q(p)=1G(p)q(p)=1-G(p) is more elastic as $p$ increases.
The absolute value of the price elasticity is $\eta(p)=-\frac{dq(p)}{dp}\cdot\frac{p}{q}=\frac{pg(p)}{1-G(p)}$. By log-concavity of 1G(p)1-G(p), $\frac{g(p)}{1-G(p)}$ is positive and increasing. Thus, $\eta(p)$ is positive and increasing for $p>0$. ∎
Higher prices therefore place the monopolist on a more price-sensitive segment of demand. The profit markup, expressible as the semi-elasticity term $qP^{\prime}(q)$, shrinks as the firm moves into more elastic regions. When cost $c$ rises, the marginal-revenue curve intersects cost at a lower quantity—precisely where the demand curve is more elastic—resulting in a higher price but a smaller markup (c.f. Figure 1).
Market power, elasticity, and information rent. Price markup is commonly used to measure market power: lower elasticity implies higher markup. Yet elasticity also mirrors the information environment. A less elastic (more inelastic) demand curve corresponds to a larger pool of high-value and relatively homogeneous consumers, which in turn implies a larger information rent for the indifferent consumer. Paradoxically, a higher information rent grants market power for the monopolist, even though in mechanism design it represents “difficulty” for the designer. The appropriate interpretation of this difficulty is thus the reluctance of supplying more good (and lowering price). Therefore, demand elasticity and information rent are two sides of the same coin. Both summarize the monopolist’s ability to extract surplus and thus provide dual interpretations of market power.
4 Related Literature
Log-concave functions and distributions have been studied extensively in probability, convex geometry, and stochastic programming. Foundational work by Prekopa1973 introduced logarithmically concave measures and established key preservation properties (e.g., under marginals and integrals) that underpin modern treatments of log-concavity and the Prékopa–Leindler inequality. More systematic probability-theoretic accounts of unimodality and convexity, including log-concavity and its links to reliability theory and monotone hazard rates, are provided in DharmadhikariJoagDev1988 and surveyed in the statistics literature by SaumardWellner2014.
Within economics, An1998 gives a complete characterization of log-concavity (and log-convexity) in terms of comparative statics and stochastic dominance, clarifying its implications for single-crossing and regularity assumptions. The review by BagBerg05Logconcave synthesizes many of these ideas for log-concave probabilities and highlights applications in appraisal design in lemon markets. The properties documented here—preservation of log-concavity under integration, truncation, and economically natural transformations—are closely related to conditions used in mechanism design and contract theory; see, for example, borgers2015 for mechanism design, where monotone hazard rate plays a technical role; or BoltonConcract for contract theory, where MLRP helps model the monitoring technology.
Appendix A Omitted Proofs
A.1 Proof of Proposition 2
Proof of 1. Suppose $f$ is a nonnegative concave function. By definition, for any $x,y$ in the domain of $f$, any $\alpha\in[0,1]$,
$f(\alpha x+(1-\alpha)y)\geq\alpha f(x)+(1-\alpha)f(y).$
Taking logarithms on both sides, we get
$\log f(\alpha x+(1-\alpha)y)\geq\log(\alpha f(x)+(1-\alpha)f(y))\geq\alpha\log f(x)+(1-\alpha)\log f(y),$
where the last inequality is implied by concavity of $\log(\cdot)$. Therefore, $f$ is log-concave.
Proof of 2. Suppose $f$ is log-concave, we have, for any $x,y$ in the domain of $f$ and any $\alpha\in[0,1]$,
$\log f(\alpha x+(1-\alpha)y)\geq\alpha\log f(x)+(1-\alpha)\log f(y),$
the right-hand side of which is greater than $\min\{\log f(x),\log f(y)\}$. Therefore $\log f$ is quasi-concave. Note that the upper level sets of $f$, defined as $\{x|f(x)\geq a\}$ for $a>0$, are equivalent to $\{x|\log f(x)\geq\log a\}$, i.e., the upper level sets of $\log f$. Since $\log f$ is quasi-concave, its upper level sets are convex. Therefore, the upper level sets of $f$ are convex, confirming that $f$ is quasi-concave.
A.2 Proof of Proposition 4
The proof needs to be done in both directions.
$f$ is log-concave $\Rightarrow$ MLRP. Suppose $\log f$ is concave. For any $\theta_{1}<\theta_{2}$ and $x<x^{\prime}$, the MLRP is equivalent to
$\log f(x-\theta_{2})+\log f(x^{\prime}-\theta_{1})\leq\log f(x-\theta_{1})+\log f(x^{\prime}-\theta_{2}).$
Define $t=\frac{x^{\prime}-x}{\theta_{2}-\theta_{1}+x^{\prime}-x}\in(0,1)$, then
$x-\theta_{1}=t(x-\theta_{2})+(1-t)(x^{\prime}-\theta_{1}),\;x^{\prime}-\theta_{2}=(1-t)(x-\theta_{2})+t(x^{\prime}-\theta_{1}).$
By concavity of $\log f$,
$\log f(x-\theta_{1})\geq t\log f(x-\theta_{2})+(1-t)\log f(x^{\prime}-\theta_{1})$
and
$\log f(x^{\prime}-\theta_{2})\geq(1-t)\log f(x-\theta_{2})+t\log f(x^{\prime}-\theta_{1}).$
Adding the above two, I obtain the desired inequality corresponding to MLRP.
$f$ is log-concave $\Leftarrow$ MLRP. Suppose the density family has MLRP. For any $a<b$, set $x-\theta_{2}=a$, $x^{\prime}-\theta_{1}=b$, and $x-\theta_{1}=x^{\prime}-\theta_{2}=\frac{a+b}{2}$ (i.e., take $t=\frac{1}{2}$). The MLRP inequality then simplifies to
$\log f(a)+\log f(b)\leq 2\log f\left(\frac{a+b}{2}\right).$
This is true for any $a<b$. Therefore, $\log f$ is concave, i.e., $f$ is log-concave.
A.3 Proof of Proposition 5
Note that
$\Gamma^{\prime}(x)=\frac{\varphi^{2}-\varphi^{\prime}\Phi}{\varphi^{2}}=1+x\frac{\Phi}{\varphi}=1+x\Gamma(x)$
where the second equality comes from $\varphi^{\prime}(x)=-x\varphi(x)$. Therefore,
$\Gamma^{\prime\prime}(x)=\Gamma(x)+x\Gamma^{\prime}(x)=\Gamma(x)+x(1+x\Gamma(x))=x+(1+x^{2})\Gamma(x)$
I claim that $\Gamma^{\prime\prime}(x)\geq 0$ for all $x$, so that $\Gamma(x)$ is convex. Note that the claim is true when $x\geq 0$. What remains is the case of $x<0$. Let y=x>0y=-x>0. One wants to show
$y+(1+y^{2})\frac{\Phi(-y)}{\varphi(-y)}\geq 0,$
the left-hand side of which is
$y+(1+y^{2})\frac{1-\Phi(y)}{\varphi(y)}$
using symmetry of the standard normal distribution. It is then equivalent to proving
$k(y):=-y\varphi(y)+(1+y^{2})(1-\Phi(y))\geq 0$
It is clear that $\lim_{y\rightarrow\infty}k(y)=0$, following properties of $\Phi$ and $\varphi$. We claim that $k(y)$ weakly decreases in $y\in(0,\infty)$, so that $k(y)\geq 0$ for all $y>0$. To see this, note that
$k^{\prime}(y)$ $=-\varphi(y)-y\varphi^{\prime}(y)+2y(1-\Phi(y))-(1+y^{2})\varphi(y)$
$=-2\varphi(y)+2y(1-\Phi(y))<0$
because [^8] This comes from an upper bound for the “Mills ratio”: $1-\Phi(y)=\int_{y}^{\infty}\varphi(t)dt=-\int_{y}^{\infty}t^{-1}d\varphi(t)=\frac{\varphi(y)}{y}-\int_{y}^{\infty}\frac{\varphi(t)}{t^{2}}dt<\frac{\varphi(y)}{y}.$
$1-\Phi(y)<\frac{\varphi(y)}{y},\forall y>0.$

### v2

@@
 I summarize BagBerg05Logconcave’s review on log-concave functions, make several corrections, and augment the discussion with further results that can be useful in obtaining monotone hazard rate.
-I also provide an application of monopoly pricing.
+I also provide an application of monopoly pricing, where strict log-concavity of demand curve implies strict concavity of the revenue function.
 1 Property of Log-Concave Functions
@@
 Each consumer’s value of the unit of demand is $v\in[0,1]$ drawn i.i.d.
-from a distribution $G$ with strictly log-concave density gg.
+from a distribution $G$ with log-concave density gg.
 [^5] A weaker condition is to assume 1G(v)1-G(v) is strictly log-concave, with which all discussion in this section goes through.
@@
 The fact that the price markup decreases in cost $c$ is a distinctive feature of log-concave demand.
-This property does not hold for general demand systems with concave revenue; for example, constant-elasticity demand exhibits increasing markup.
+This property does not hold for general demand systems with concave revenue; for example, constant-elasticity demand (with elasticity larger than $1$) exhibits increasing markup.
 The intuition is most transparent from the perspective of information rents. For a log-concave distribution $G$, the information rent of type $v$, given by
@@
 Paradoxically, a higher information rent grants market power for the monopolist, even though in mechanism design it represents “difficulty” for the designer.
-The appropriate interpretation of this difficulty is thus the reluctance of supplying more good (and lowering price).
+The appropriate interpretation of this difficulty is thus the reluctance of supplying more good (and lowering price) when making the marginal tradeoff.
 Therefore, demand elasticity and information rent are two sides of the same coin.